OpenStudy (anonymous):

write an equation of the line that passes through (-5,8) and (-3,-8)

OpenStudy (anonymous):

slope = [8-(-8)]/[-5-(-3)] =16/-2 = -8 y=mx+b 8=-8(-5)+b 8=40+b b=-32 y=-8x-32

OpenStudy (anonymous):

\[m=\frac{y1-y2}{x1-x2}=\frac{8+8}{-5+3}=-8\]